# A Survey of LLM-based Automated Program Repair: Taxonomies, Design Paradigms, and Applications

### v1

@@ 1. Introduction
Table 2. Annual distribution of LLM-based APR applications by paradigm
@@
2023 7 6 1 0 14
2024 10 4 9 7 30
2025 (Jan - June) 4 2 4 7 17
Total 21 14 14 14 63
Our survey distills several insights into the trade-offs among the four LLM-based APR paradigms. Fine-tuning aligns a model tightly with the repair task, producing highly accurate fixes for defects similar to those in its training data; however, it demands large bug-fix corpora and considerable GPU time. Prompting requires no additional training and supports rapid deployment, yet its effectiveness hinges on careful prompt engineering and is bounded by the model’s fixed knowledge and context window. Both paradigms can be strengthened with retrieval-augmented or analysis-augmented prompts that inject external knowledge or program-analysis feedback, thereby improving semantic correctness and reducing invalid patches. Procedural pipelines embed one or more LLM calls in deterministic multi-step workflows, offering reproducible control at moderate runtime cost. Agentic frameworks delegate planning and tool invocation to the LLM itself, enabling multi-hunk and cross-file repair at the price of increased engineering complexity, latency, and resource consumption. This community shift from single-turn generation to iterative, tool-assisted workflows reflects a drive to boost success rates on complex defects. Nonetheless, three challenges remain: verifying semantic correctness beyond limited test suites, repairing bugs that span multiple modules or exceed context limits, and mitigating the substantial computational cost of large-model pipelines. Addressing these issues is crucial for making LLM-based APR reliable and efficient in continuous integration settings.
Our contributions.
  1. (i)
    We present the first unified taxonomy that covers four repair paradigms and two augmentation layers for LLM-based APR.
  2. (ii)
    We review 63 recent state-of-the-art APR systems published between Jan 2022 and June 2025, summarizing their design trade-offs, common trends, and open questions.
  3. (iii)
    We maintain a living bibliography at https://github.com/GLEAM-Lab/ProgramRepair.
@@ 2. Fine-Tuning Approaches
Fine-tuning adapts a pre-trained LLM for APR by updating its weights using bug-fix data. Research in this area now revolves around five distinct strategies. Fully supervised fine-tuning updates every parameter of the backbone and achieves strong accuracy across diverse benchmarks (Jiang et al., 2023a; Zhou et al., 2024b; Jiang and Wu, 2024). Parameter-efficient fine-tuning freezes most of the backbone and learns small adapter modules, which lowers memory cost while keeping competitive repair performance (Silva et al., 2023; Yang et al., 2025b; Luo et al., 2024; Li et al., 2024c; Ruiz et al., 2025). Reinforcement learning fine-tuning trains the model with execution or security rewards so that it optimizes for functional correctness instead of log likelihood (Zhao et al., 2024a; Islam et al., 2024; Wei et al., 2025; Dai et al., 2025; Yu et al., 2025a). Knowledge distillation transfers bug-fixing skill from a large teacher or rule set to a smaller student, cutting inference latency without much accuracy loss (Jiang et al., 2023b; Wong et al., 2024; Yang et al., 2024d). Context-enriched fine-tuning adds static-analysis results, execution traces, or retrieved patches to the input so the model can reason beyond the local snippet (Bouzenia et al., 2023; Jin et al., 2023; Chow et al., 2024; Huang et al., 2024b).
@@ 2.1. Full Fine-Tuning
With the emergence of code-target language models based on transformers (Vaswani et al., 2017) such as GPT-2 (Radford et al., 2019), CodeT5 (Wang et al., 2021), and CodeLlama (Roziere et al., 2023), researchers have pursued full-model supervised fine-tuning on program repair scenarios. Recent work explores several variants of this strategy. Large-scale benchmark adaptation fine-tunes multiple backbone LLMs across standard datasets to quantify the effects of size and data (Jiang et al., 2023a). Security-specific pipelines inject vulnerability patterns and CWE knowledge during full tuning to target secure patching (Zhou et al., 2024b). Domain-focused fine-tuning demonstrates that even 1B-parameter LLMs can learn high-yield repair heuristics when specialized data are available (Jiang and Wu, 2024). These efforts illustrate the breadth of full-parameter supervised fine-tuning for APR.
Huang et al. conducted an empirical study on fine-tuning LLMs (including CodeBERT (Feng et al., 2020), GraphCodeBERT (Guo et al., 2021), PLBART (Ahmad et al., 2021), CodeT5 (Wang et al., 2021), and UniXcoder (Guo et al., 2022)) for program repair tasks (Huang et al., 2023a). Experimental results show that LLMs in the fine-tuning paradigm can significantly outperform previous state-of-the-art APR tools on both single-hunk and multi-hunk bugs within Defects4J (Just et al., 2014). However, they reveal that LLMs usually fail to generate correct patches due to the input/output length limit of LLMs. On the other hand, the generated candidate patches are significantly limited, as hardware resources often fail to meet the high computational demands of LLMs for massive candidate patches.
Jiang et al. conducted a large-scale study that fully fine-tunes ten pre-trained code language models across four APR benchmarks to quantify the effect of supervised adaptation on repair performance (Jiang et al., 2023a). Fine-tuning raised fix rates by up to an order of magnitude and showed a clear positive correlation between model size and patch success. The work also revealed heavy GPU demand and frequent overfitting to the suspected bug location, highlighting concerns about efficiency and generalization.
@@ 2.2. Parameter-Efficient Fine-Tuning
Parameter-efficient fine-tuning (PEFT) alleviates the memory and compute burden of adapting LLMs for program repair by updating only a compact set of task-specific parameters. Current PEFT strategies for repair fall into three families: (i) structural adapters, including adapter layers (Houlsby et al., 2019), prefix vectors (Lester et al., 2021), and low-rank matrices such as LoRA (Hu et al., 2021); (ii) quantization-aware adapters typified by QLoRA, which merges 4-bit quantization with LoRA (Dettmers et al., 2023); and (iii) noise-regularized tuning exemplified by NEFTune, which injects calibrated embedding noise (Jain et al., 2023a). These designs shrink trainable parameters by up to three orders of magnitude, enable single-GPU training, and keep the frozen backbone reusable for other tasks. Efficiency-oriented variants demonstrate promising effectiveness in program and vulnerability repair, broadening PEFT’s applicability to larger models and consumer-grade hardware (Silva et al., 2023; Yang et al., 2025b; Luo et al., 2024). Nevertheless, attaining strong repair accuracy still requires judicious adapter placement, appropriate rank selection, and high-quality bug-fix data, because limited adapter capacity can miss semantic links that span multiple files (Li et al., 2024c; Ruiz et al., 2025).
RepairLLaMA augments a frozen CodeLlama-7B with rank-8 LoRA matrices (0.06% of parameters) and fine-tunes them on 4,000 curated bug-fix pairs (Silva et al., 2023). Training on a single GPU fixes 144 Defects4J-V2 (Just et al., 2014) bugs, 109 HumanEval-Java tasks (Jiang et al., 2023a), and 20 GitBug-Java issues (Silva et al., 2024), outperforming larger frozen models. Limited adapter capacity and reliance on static fault localization reduce accuracy for multi-hunk or cross-file defects.
@@
Luo et al. apply LoRA adapters in a federated setting where each organization trains locally on proprietary code and a server aggregates updates, then evaluates on EvalRepair-Java (Luo et al., 2024). The protocol preserves privacy and matches centralized training within two percentage points while exchanging only megabytes of adapter weights per round. Convergence slows under highly heterogeneous client data, and a secure aggregation infrastructure is required.
Li et al. benchmark LoRA, prefix-tuning, adapters, and (IA)3 (Liu et al., 2022) on the APR-INSTRUCTION corpus with CodeT5+ and StarCoder backbones (Li et al., 2024c). The strongest adapter (IA)3 resolves 58% more bugs than earlier LLM baselines while keeping memory usage low, but its gains plateau once adapter size or data volume exceeds moderate thresholds, so careful hyper-parameter tuning remains essential.
@@
2.3. Knowledge Distilling
Knowledge distillation transfers bug-fixing skills from a high-capacity teacher or knowledge source to a smaller student, allowing inference costs to drop without compromising accuracy. Recent studies have applied classical teacher-student distillation (including soft-target and rationale distillation) (Wong et al., 2024), dataset distillation (Yang et al., 2024d), and rule-guided distillation (Jiang et al., 2023b) to APR, yielding strong empirical gains. These techniques either compress large teachers or replace them with structured knowledge, inject structural hints, restrict generation using structured rules, or enforce domain constraints while keeping the student lightweight. The following paragraphs review three representative methods that illustrate the breadth of distillation for program repair.
KNOD distills formal syntax and semantic rules into a student decoder by aligning its token probabilities with a rule-derived teacher distribution (Jiang et al., 2023b). The method repaired 72 Defects4J v1.2 bugs, 25 QuixBugs, and 50 Defects4J v2.0 bugs, establishing new benchmark records. By restricting token choices to grammar-compliant options, KNOD greatly improves compile success and patch validity. Rule coverage is limited to statically checkable properties and offers less help for logical faults that do not violate syntax or types.
@@ 2.3. Knowledge Distilling
NARRepair is a non-autoregressive APR model that first distills a CodeT5-large (Wang et al., 2021) teacher to build a cleaner student training set, then generates patches in parallel using action guidance and AST dependencies (Yang et al., 2024d). It repairs 69, 45, and 23 bugs on Defects4J (Just et al., 2014) v1.2, v2.0, and QuixBugs (Lin et al., 2017) while delivering 5.4-18.6 times faster inference than autoregressive baselines. The combination of dataset distillation and two-stage decoding bridges the usual quality gap of non-autoregressive generation without sacrificing speed. However, all evaluations are Java-only, so cross-language generality is unverified.
@@ 2.4. Reinforcement Learning Fine-Tuning
Recently, policy-gradient algorithms have been widely used in LLM-based research, including PPO (Schulman et al., 2017), DAPO (Yu et al., 2025b), and Group Relative Policy Optimization GRPO (Shao et al., 2024). In this section, we also include preference-based tuning, such as DPO (Rafailov et al., 2023), which is not strictly an RL algorithm. Recent RL-based systems include RePair (Zhao et al., 2024a), SecRepair (Islam et al., 2024), and SWE-RL (Wei et al., 2025). A preference-based alternative, AdaPatcher (Dai et al., 2025) and Smart-LLaMA-DPO (Yu et al., 2025a), adopt Direct Preference Optimization (DPO) to refine their patch policy. In program-repair settings, the reward most often combines execution signals, such as compile-and-test success, with static cues like vulnerability removal or diff similarity to reduce sparsity (Shojaee et al., 2023; Ye et al., 2025). Some recent works adopt a two-stage pipeline that performs supervised fine-tuning to obtain a stable repair model and then applies RLFT for further alignment (Dai et al., 2025). Compared with purely supervised fine-tuning, RLFT can push the model beyond the training distribution and align it directly with functional correctness; however, it incurs heavy computation because each gradient update may require running tests, rewards are often sparse, and poorly shaped rewards can lead to over-fitting weak test oracles (Dou et al., 2024; Liu et al., 2023; Ye et al., 2025).
RePair starts from a fully supervised StarCoder-15B model and adds an iterative RL loop miming human debugging (Zhao et al., 2024a). A smaller critic model, trained to distinguish good from bad patches based on compile and test outcomes, supplies rewards to the actor via PPO. The actor may propose several successive fixes for one bug, receiving intermediate feedback at each step. This process enhances repair success on competitive-programming bugs to a level comparable to that of much larger proprietary LLMs. The trade-off is higher latency and the need to guard against infinite repair loops on hard-to-fix defects.
SecRepair targets vulnerability repair. It slices the code around each flaw, then fine-tunes a CodeGen2 (Nijkamp et al., 2023) actor with PPO where the reward combines unit-test outcomes, semantic-similarity checks, and a security bonus that favors patches eliminating unsafe patterns (Islam et al., 2024). Trained on their built 36k-example InstructVul dataset and evaluated on zero- and N-day bugs from 6 open-source IoT operating systems, SecRepair fixes 58% of vulnerable functions and generates concise commit messages that describe the flaw, outperforming a supervised baseline by up to 22 percentage points F1 while adding only 9% runtime overhead. Strengths include security-aware rewards and explicit vulnerability descriptions; limitations are reliance on known flaw locations and increased token usage for descriptive comments.
@@
AdaPatcher adopts a strict two‑stage pipeline: it first performs supervised fine‑tuning on CodeLlama‑Instruct‑7B to train a Bug Locator that pinpoints faulty lines, then applies reinforcement learning, in the form of Direct Preference Optimization (DPO) (Rafailov et al., 2023), to a Program Modifier that refines the patch policy using preference pairs derived from unit‑test outcomes and a “fewer‑changes” reward, achieving 62% consistency and 67.57% accuracy on their newly released ACPR benchmark, a clear lead over purely supervised or prompt‑based baselines (Dai et al., 2025). Its strength lies in marrying location‑aware supervised fine-tuning with preference‑based RLFT: the first stage supplies reliable edit anchors, and the second stage’s lightweight reward (diff similarity + compile‑and‑test bonus) lets the model learn to generate minimal yet correct edits without the high runtime of full Monte‑Carlo execution.
Smart-LLaMA-DPO targets smart-contract vulnerability detection and explanation through a three-stage post-training pipeline on a LLaMA-3.1-8B backbone (Yu et al., 2025a). First, a continual pre-training (CPT) phase feeds millions of curated Solidity contracts to solidify domain knowledge. Second, supervised fine-tuning (SFT) optimizes a dual-objective loss that balances detection logits and explanation likelihoods on a hand-verified dataset covering four major vulnerability types and seven machine-unauditable (MU) categories. Finally, the model undergoes Direct Preference Optimization (DPO) (Rafailov et al., 2023), a lightweight, reward-free RL method that updates policy logits to favor expert-preferred explanations over suboptimal ones, using paired outputs crafted by security specialists. On the same evaluation suite, Smart-LLaMA-DPO raises F1 by 7.5%-11.1% and accuracy by up to 10.5% over the previous SOTA DMT (Qian et al., 2023), and beats iAudit (Ma et al., 2024) on seven MU vulnerability classes by +26% F1 and +18% accuracy, while human judges mark 81%, 84%, and 95% of its explanations as correct, thorough, and clear, respectively. However, its current scope is limited to Solidity and single-file contracts, and scaling the preference dataset remains an open challenge.

### v2

@@ 3.4. Annual Adoption by Paradigm
Table 4. Annual distribution of LLM-based software repair applications by paradigm
@@
2023 5 6 1 0 12
2024 9 2 5 3 19
2025 7 4 8 10 29
Total 21 14 14 13 62
Answer to RQ1 In our select works, adoption shifts from single-turn prompting toward procedural and agentic workflows between 2022 and October 2025. We observe growing use of retrieval and analysis augmentation across all paradigms, while fine-tuning remains common when high task alignment is required.
@@ 4. Fine-Tuning Approaches
Fine-tuning adapts a pre-trained LLM for software repair by updating its weights using bug-fix data. Research in this area now revolves around five distinct strategies. Fully supervised fine-tuning updates every parameter of the backbone and achieves strong accuracy across diverse benchmarks (Jiang et al., 2023a; Zhou et al., 2024b; Jiang and Wu, 2024). Parameter-efficient fine-tuning freezes most of the backbone and learns small adapter modules, which lowers memory cost while keeping competitive repair performance (Silva et al., 2025; Yang et al., 2025b; Luo et al., 2025b; Li et al., 2024c; Ruiz et al., 2025). Reinforcement learning fine-tuning trains the model with execution or security rewards so that it optimizes for functional correctness instead of log likelihood (Zhao et al., 2024a; Islam et al., 2024; Wei et al., 2025; Dai et al., 2025; Yu et al., 2025a). Knowledge distillation transfers bug-fixing skill from a large teacher or rule set to a smaller student, cutting inference latency without much accuracy loss (Jiang et al., 2023b; Wong et al., 2024; Yang et al., 2024c). Context-enriched fine-tuning injects static-analysis results, execution traces, or retrieved patches into the training input (Bouzenia et al., 2023; Jin et al., 2023; Chow et al., 2024; Huang et al., 2024b).
@@ 4.1. Full Fine-Tuning
With the emergence of code-target language models based on transformers (Vaswani et al., 2017) such as DeepSeek-Coder (Guo et al., 2024; Zhu et al., 2024), CodeT5 (Wang et al., 2021), and CodeLlama (Roziere et al., 2023), researchers have pursued full-model supervised fine-tuning on program repair scenarios. Recent work explores several variants of this strategy. Large-scale benchmark adaptation fine-tunes multiple backbone LLMs across standard datasets to quantify the effects of size and data (Jiang et al., 2023a). Security-specific pipelines inject vulnerability patterns and CWE knowledge during full tuning to target secure patching (Zhou et al., 2024b). Domain-focused fine-tuning demonstrates that even 1B-parameter LLMs can learn high-yield repair heuristics when specialized data are available (Jiang and Wu, 2024). These efforts illustrate the breadth of full-parameter supervised fine-tuning for software repair.
Huang et al. conducted an empirical study on fine-tuning LLMs (including CodeBERT (Feng et al., 2020), GraphCodeBERT (Guo et al., 2021), PLBART (Ahmad et al., 2021), CodeT5 (Wang et al., 2021), and UniXcoder (Guo et al., 2022)) for program repair tasks (Huang et al., 2023a). Experimental results show that LLMs in the fine-tuning paradigm can significantly outperform previous state-of-the-art repair tools on both single-hunk and multi-hunk bugs within Defects4J (Just et al., 2014). However, they reveal that LLMs usually fail to generate correct patches due to the input/output length limit of LLMs. On the other hand, the generated candidate patches are significantly limited, as hardware resources often fail to meet the high computational demands of LLMs for massive candidate patches.
Jiang et al. conducted a large-scale study that fully fine-tunes ten pre-trained code language models across four repair benchmarks to quantify the effect of supervised adaptation on repair performance (Jiang et al., 2023a). Fine-tuning raised fix rates by up to an order of magnitude and showed a clear positive correlation between model size and patch success. The work also revealed heavy GPU demand and frequent overfitting to the suspected bug location, highlighting concerns about efficiency and generalization.
@@ 4.2. Parameter-Efficient Fine-Tuning
Parameter-efficient fine-tuning (PEFT) alleviates the memory and compute burden of adapting LLMs for program repair by updating only a compact set of task-specific parameters. Current PEFT strategies for repair fall into three families: (i) structural adapters, including adapter layers (Houlsby et al., 2019), prefix vectors (Lester et al., 2021), and low-rank matrices such as LoRA (Hu et al., 2021); (ii) quantization-aware adapters typified by QLoRA, which merges 4-bit quantization with LoRA (Dettmers et al., 2023); and (iii) noise-regularized tuning exemplified by NEFTune, which injects calibrated embedding noise (Jain et al., 2023a). These designs shrink trainable parameters by up to three orders of magnitude, enable single-GPU training, and keep the frozen backbone reusable for other tasks. Efficiency-oriented variants demonstrate promising effectiveness in program and vulnerability repair, broadening PEFT’s applicability to larger models and consumer-grade hardware (Silva et al., 2025; Yang et al., 2025b; Luo et al., 2025b). Nevertheless, attaining strong repair accuracy still requires judicious adapter placement, appropriate rank selection, and high-quality bug-fix data, because limited adapter capacity can miss semantic links that span multiple files (Li et al., 2024c; Ruiz et al., 2025).
RepairLLaMA augments a frozen CodeLlama-7B with rank-8 LoRA matrices (0.06% of parameters) and fine-tunes them on 4,000 curated bug-fix pairs (Silva et al., 2025). Training on a single GPU fixes 144 Defects4J v2 (Just et al., 2014) bugs, 109 HumanEval-Java tasks (Jiang et al., 2023a), and 20 GitBug-Java issues (Silva et al., 2024), outperforming larger frozen models. Limited adapter capacity and reliance on static fault localization reduce accuracy for multi-hunk or cross-file defects.
@@
Luo et al. apply LoRA adapters in a federated setting where each organization trains locally on proprietary code and a server aggregates updates, then evaluates on EvalRepair-Java (Luo et al., 2025b). The protocol preserves privacy and matches centralized training within two percentage points while exchanging only megabytes of adapter weights per round. Convergence slows under highly heterogeneous client data, and a secure aggregation infrastructure is required.
Li et al. benchmark LoRA, prefix-tuning, adapters, and (IA)3 (Liu et al., 2022) with CodeT5+ and StarCoder backbones (Li et al., 2024c). The strongest adapter (IA)3 resolves 58% more bugs than earlier LLM baselines while keeping memory usage low, but its gains plateau once adapter size or data volume exceeds moderate thresholds, so careful hyper-parameter tuning remains essential.
@@
4.3. Knowledge Distillation
Knowledge distillation transfers bug-fixing skills from a high-capacity teacher or knowledge source to a smaller student, allowing inference costs to drop without compromising accuracy. Recent studies have applied classical teacher-student distillation (including soft-target and rationale distillation) (Wong et al., 2024), dataset distillation (Yang et al., 2024c), and rule-guided distillation (Jiang et al., 2023b) to software repair, yielding strong empirical gains. These techniques either compress large teachers or replace them with structured knowledge, inject structural hints, restrict generation using structured rules, or enforce domain constraints while keeping the student lightweight. The following paragraphs review three representative methods that illustrate the breadth of distillation for program repair.
KNOD distills formal syntax and semantic rules into a student decoder by aligning its token probabilities with a rule-derived teacher distribution (Jiang et al., 2023b). The method repaired 72 Defects4J v1.2 bugs and 50 Defects4J v2.0 bugs, establishing new benchmark records. By restricting token choices to grammar-compliant options, KNOD greatly improves compile success and patch validity. Rule coverage is limited to statically checkable properties and offers less help for logical faults that do not violate syntax or types.
@@ 4.3. Knowledge Distillation
NARRepair first distills a CodeT5-large (Wang et al., 2021) teacher to build a cleaner student training set, then generates patches in parallel using action guidance and AST dependencies (Yang et al., 2024c). It repairs 69, 41, and 23 bugs on Defects4J (Just et al., 2014) v1.2, v2.0, and QuixBugs (Lin et al., 2017) while delivering 5.4-18.6 times faster inference than autoregressive baselines. The combination of dataset distillation and two-stage decoding bridges the usual quality gap of non-autoregressive generation without sacrificing speed. However, all evaluations are Java-only, so cross-language generality is unverified.
@@ 4.4. Reinforcement Learning Fine-Tuning
Recently, policy-gradient algorithms have been widely used in LLM-based research, including PPO (Schulman et al., 2017), DAPO (Yu et al., 2025b), and Group Relative Policy Optimization GRPO (Shao et al., 2024). In this section, we also include preference-based tuning, such as DPO (Rafailov et al., 2023), which is not strictly an RL algorithm. Recent RL-based systems include RePair (Zhao et al., 2024a), SecRepair (Islam et al., 2024), SWE-RL (Wei et al., 2025), and Vul-R2 (Wen et al., 2025), while a preference-based alternative, AdaPatcher (Dai et al., 2025), adopts Direct Preference Optimization (DPO) to refine its patch policy. In program-repair settings, the reward most often combines execution signals, such as compile-and-test success, with static cues like vulnerability removal or diff similarity to reduce sparsity (Shojaee et al., 2023; Ye et al., 2025). Some recent works adopt a two-stage pipeline that performs supervised fine-tuning to obtain a stable repair model and then applies RLFT for further alignment (Dai et al., 2025). Compared with purely supervised fine-tuning, RLFT can push the model beyond the training distribution and align it directly with functional correctness; however, it incurs heavy computation because each gradient update may require running tests, rewards are often sparse, and poorly shaped rewards can lead to over-fitting weak test oracles (Dou et al., 2024; Liu et al., 2023; Ye et al., 2025).
RePair starts from a fully supervised StarCoder-15B model and adds an iterative RL loop mimicking human debugging (Zhao et al., 2024a). A smaller critic model, trained to distinguish good from bad patches based on compile and test outcomes, supplies rewards to the actor via PPO. The actor may propose several successive fixes for one bug, receiving intermediate feedback at each step. This process enhances repair success on competitive-programming bugs to a level comparable to that of much larger proprietary LLMs. The trade-off is higher latency and the need to guard against infinite repair loops on hard-to-fix defects.
SecRepair targets vulnerability repair. It slices the code around each flaw, then fine-tunes a CodeGen2 (Nijkamp et al., 2023) actor with PPO, where the reward combines unit-test outcomes, semantic-similarity checks, and a security bonus that favors patches eliminating unsafe patterns (Islam et al., 2024). Trained on their built 36k-example InstructVul dataset and evaluated on zero- and N-day bugs from 6 open-source IoT operating systems, SecRepair fixes 58% of vulnerable functions and generates concise commit messages that describe the flaw, outperforming a supervised baseline by up to 22 percentage points F1 while adding only 9% runtime overhead. Strengths include security-aware rewards and explicit vulnerability descriptions; limitations are reliance on known flaw locations and increased token usage for descriptive comments.
@@
AdaPatcher adopts a strict two-stage pipeline: it first performs supervised fine-tuning on CodeLlama-Instruct-7B to train a Bug Locator that pinpoints faulty lines, then applies reinforcement learning, in the form of Direct Preference Optimization (DPO) (Rafailov et al., 2023), to a Program Modifier that refines the patch policy using preference pairs derived from unit-test outcomes and a “fewer-changes” reward, achieving 62% consistency and 67.57% accuracy on their newly released ACPR benchmark, a clear lead over purely supervised or prompt-based baselines (Dai et al., 2025). Its strength lies in marrying location-aware supervised fine-tuning with preference-based RLFT: the first stage supplies reliable edit anchors, and the second stage’s lightweight reward (diff similarity + compile-and-test bonus) lets the model learn to generate minimal yet correct edits without the high runtime of full Monte-Carlo execution.
Vul-R2 treats vulnerability repair as a reasoning task, training a Qwen-based model that produces explanations and patches for C/C++ code through a two-stage pipeline: supervised fine-tuning for reasoning first, then reinforcement learning from verifiable rewards (Wen et al., 2025). In its Curriculum-based Verifiable Rewarded Training phase, an RLVR objective uses a critic LLM’s fix-or-not judgment plus ground-truth similarity and a simple format reward to score each patch. On the PrimeVul (Ding et al., 2025) and SVEN (He and Vechev, 2023) benchmarks, this RL phase significantly improves exact match and CodeBLEU over SFT and CoT baselines, and these gains coincide with longer, more reflective repair traces, giving unusually direct evidence that RLFT can improve reasoning rather than simply overfitting tests.